# State-Constrained Offline Reinforcement Learning

### v1

@@ Abstract
Traditional offline reinforcement learning methods predominantly operate in a batch-constrained setting. This confines the algorithms to a specific state-action distribution present in the dataset, reducing the effects of distributional shift but restricting the algorithm greatly. In this paper, we alleviate this limitation by introducing a novel framework named state-constrained offline reinforcement learning. By exclusively focusing on the dataset’s state distribution, our framework significantly enhances learning potential and reduces previous limitations. The proposed setting not only broadens the learning horizon but also improves the ability to combine different trajectories from the dataset effectively, a desirable property inherent in offline reinforcement learning. Our research is underpinned by solid theoretical findings that pave the way for subsequent advancements in this domain. Additionally, we introduce StaCQ, a deep learning algorithm that is both performance-driven on the D4RL benchmark datasets and closely aligned with our theoretical propositions. StaCQ establishes a strong baseline for forthcoming explorations in state-constrained offline reinforcement learning.
@@ 1 Introduction
Offline reinforcement learning (RL) focuses on deriving an optimal decision-making policy from a fixed dataset of previously collected experiences [36, 38]. It prohibits online exploration, meaning that effective policies must be constructed solely from the evidence provided by an unknown, potentially sub-optimal behaviour policy active in the environment. This approach is especially suited for real-world scenarios where executing sub-optimal actions can be dangerous, time-consuming, or costly, yet there is an abundance of prior data. Relevant applications include robotics [46, 34, 47], long-term healthcare treatment plans [51, 50, 44], and autonomous driving [43, 18, 16]. Despite its practical appeal, offline RL faces a significant challenge called distributional shift [33]. This phenomenon arises when attempting to estimate the values of actions not present in the dataset, often manifesting as an overestimation. As a result, out-of-distribution (OOD) actions are perceived as more valuable than they truly are, causing the agent to select sub-optimal actions and leading to the accumulation of errors [22].
Currently available offline RL techniques mitigate the issue of distributional shift through two primary approaches. The first confines the policy to produce only actions likely to emerge from the dataset or its distribution [22, 33, 55, 45, 31, 63, 21]. The second approach revolves around conservatively estimating the value of OOD actions [35, 58, 3]. Both strategies aim to anchor the learning algorithm to the state-action pairs within the dataset, an approach termed batch-constrained [22]. By doing so, these methods seek to minimise the impact of distributional shift. However, adhering strictly to the state-action distribution can be restrictive. Thus, recent techniques seek a balance between exploiting potentially valuable OOD actions and minimising distributional shift repercussions.
@@
Constraining exclusively to the dataset’s state distribution could significantly reduce the requisite dataset size. Instead of needing all state-action pairs, the focus shifts to states alone. This approach allows the agent to initiate out-of-distribution (OOD) actions, provided they lead to a known state. Such flexibility can enhance the capability of offline RL algorithms in trajectory stitching—combining sub-optimal experiences into improved trajectories. Rather than connecting based on actions from different trajectories, the state-constrained approach facilitates stitching by leveraging adjacent states. This method relies on understanding these proximate states, a concept we label as reachability. This concept refers to the ability to reach a particular state from another state within the dataset. In the state-constrained approach, understanding reachability is crucial for effectively stitching together sub-optimal trajectories and improving the overall policy. In some environments, reachability is easily delineated; for example, in grid-based environments like mazes, where the agent can move directly between adjacent cells. In other environments, reachability must be ascertained from the dataset, such as in complex robotic systems where state transitions depend on intricate dynamics.
@@
Building upon the concept of reachability, in this paper, we develop a state-constrained offline RL approach and make several contributions to this new area. In the batch-constrained approach, the learning algorithm is anchored to the state-action pairs within the dataset. In contrast, the state-constrained approach focuses solely on the states, allowing for more flexibility in action selection while still mitigating distributional shift. We provide our state-constrained framework with theoretical guarantees, built on a few foundational assumptions. We demonstrate that the state-constrained framework will always construct a policy that produces higher or equal-value actions compared to the batch-constrained policy. Finally, we propose and empirically validate a novel deep state-constrained offline RL algorithm, called StaCQ. StaCQ learns a state-constrained Q-function and updates the policy to remain close to the highest quality reachable states within the dataset. Our experiments show that StaCQ demonstrates competitive performance, outperforming several state-of-the-art methods on the D4RL benchmarking dataset [20] in many tasks. StaCQ outperforms all model-free baselines in the locomotion tasks and shows strong performance in the Antmaze tasks. These results provide a strong basis for many future state-constrained algorithms to be built upon. We believe that this work will serve as a foundation for the development of future state-constrained algorithms, much like how BCQ [22] has done for batch-constrained techniques, and contribute to the advancement of offline RL.
2 Related work
Model-free offline RL.
BCQ [22] emerged as one of the pioneer offline deep RL methodologies. It posits that possessing all $(s,a)$-pairs in the MDP allows for constructing an optimal QA-value by narrowing the update solely to $(s,a)$-pairs present in the dataset. Consequently, BCQ restricts the policy to $(s,a)$-pairs in the dataset while allowing minimal perturbations to facilitate marginal benefits. Post-BCQ, numerous algorithms have been proposed to devise alternative means to limit the policy to $(s,a)$-pairs in the dataset, be it in the support of the $(s,a)$-distribution [33, 55, 45, 31, 63, 8] or directly to the $(s,a)$-pairs [21]. In a distinct vein, instead of policy constraints, certain methods have sought to circumvent distributional shifts by pessimistically evaluating the value function on out-of-distribution $(s,a)$-pairs [35, 58, 3]. A subset of offline RL methodologies both pessimistically evaluate the value function and constrain the policy [14, 7]. Both clusters of methods adopt the batch-constrained approach which prioritises $(s,a)$-pairs in the dataset over unobserved actions. Our state-constrained method eases the batch-constrained objective, necessitating constraint only to dataset states rather than $(s,a)$-pairs.
Model-based offline RL.
Model-based methods learn the dynamics of the environment to aid policy learning [48, 28]. There is usually two main approaches to model-based offline RL. The first approach learns a pessimistic model of the environment and performs rollouts using this augmented model, increasing the dataset size [59, 29, 42]. The other main approach aims to use the models to perform planning [4, 60, 27, 15]. This means that, during evaluation the models are used to look in the future to decide the optimal path for the agent. In contrast, StaCQ uses the models to implement our notion of state reachability. This means that the models are not used to imagine new states, and there is no planning involved at evaluation.
State reachability.
The state-constrained strategy heavily leans on state reachability. Earlier studies have explored state reachability [25, 26], with a Gaussian distribution determining the stitching feasibility of $s^{\prime}$ from $s$. Our approach, in contrast, deems a state as either reachable or not, steering clear of continuous probability distributions. The notion of state reachability aligns with state similarity, a subject of extensive study in literature [61, 1, 37]. One technique to gauge state similarity is through bisimulation metrics grounded in the environment’s dynamics [19]. Bisimulations usually mandate full state enumeration [11, 5, 6], prompting the introduction of a scalable pseudometric approach [10]. Integrating a pseudometric into our state reachability concept would be intriguing, but such exploration is deferred to future research. Our metric for state reachability is intuitively graspable and stems directly from the foundational definition.
Trajectory stitching.
Offline RL’s hallmark is its stitching capability [32]. This entails amalgamating previously observed trajectories to produce a novel trajectory that accomplishes the task. Imitation learning strategies, such as behavioural cloning (BC) [40, 41], are ineffective at stitching, given their inability to discern between optimal and sub-optimal states. Analogous to BC, the Decision transformer (DT)[12] converts the offline RL challenge into a supervised learning task by employing a goal-conditioned policy. Despite DT’s deficient stitching capabilities, efforts have been undertaken to rectify this by incorporating offline RL characteristics [57, 56]. Integrating DT with the policy extraction method of StaCQ may offer potential insights into enhancing stitching capabilities. However, our present study is squarely centred on a basic implementation of state-constrained offline RL.
@@ 3 Preliminaries
The RL framework involves an agent interacting with an environment, which is typically represented as a Markov Decision Process (MDP). Such an MDP is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively; $P=p(s^{\prime}|s,a)$ represents the environment transition dynamics; $R=r(s,a,s^{\prime})$ is the reward function for transitioning; and $10\leq\gamma<$ 1 is the discount factor [49]. We focus on the deterministic MDP, where $p(s^{\prime}|s,a)=\{0,1\}$ and $r(s,a,s^{\prime})=r(s,s^{\prime})$. In RL, the agent’s objective is to identify an optimal policy, $\pi(s)$, that maximises the future discounted sum of rewards $\sum_{i=t}\gamma^{i-t}r_{i}(s_{i},s^{\prime}_{i})$.
QA-values, $Q(s,a)=\mathbb{E}_{\pi}[\sum_{i=0}\gamma^{t}r_{t}(s_{t},a_{t})|s_{0}=s,a_{0}=a]$, are the expected future discounted rewards for executing action $a$ in state $s$ and there after following policy $\pi$. Q-learning [54], denoted as QA-learning in this paper, estimates the QA-values under the assumption that future actions are chosen optimally,
@@
The optimal policy derived from QA-learning identifies the best action that maximises the QA-values, $\pi^{*}(s)=\text{arg}\max_{a}Q(s,a)$. In deterministic contexts, QA-learning parallels QS-learning [17], which estimates the expected rewards upon transitioning from state $s$ to $s^{\prime}$, followed by optimal decisions:
@@
QS-learning enables QS-values to be learned for transitioning, without evaluating actions. The optimal policy via QS-learning discerns the most advantageous subsequent state to maximise QS-values, $\pi_{s}^{*}(s)=\text{arg}\max_{s^{\prime}}Q(s,s^{\prime})$. An action can then be retrieved from an inverse dynamics model as $a=I(s,\pi_{s}^{*}(s))$, a strategy designed to address QA-learning’s challenges in redundant action spaces. Redundant action spaces refer to situations where multiple actions lead to the same next state, making it challenging for QA-learning to distinguish between them.
In offline RL, the agent aims to discover an optimal policy; however, it must solely rely on a fixed dataset without further interactions with the environment [36, 38]. The dataset comprises trajectories made up of transitions consisting of the current state, action, next state, and the reward for transitioning, where actions are chosen based on an unknown behaviour policy, $\pi_{\beta}$. Batch-Constrained QA-learning (BCQL) [22] adapts QA-learning for the offline setting by restricting the optimisation to state-action pairs present in the dataset:
@@
This approach is restrictive since convergence to the optimal value function is only ensured if every conceivable $(s,a)$-pair from the MDP resides within the dataset. Such a limitation aims to sidestep extrapolation errors resulting from distributional shifts. Nonetheless, this framework places a significant constraint on any learning algorithm. Subsequent sections relax this constraint: instead of adhering strictly to state-action pairs (batch-constrained), the learning process is only bound by states (state-constrained). In the following section, we introduce the state-constrained framework and provide a new algorithm called state-constrained QS-learning (SCQL). We show that, under minor assumptions, SCQL converges to the optimal QS-value and produces a less-restrictive policy than BCQL.
4 State-constrained QS-learning
In this section, we provide a formal introduction to state-constrained QS-learning and establish its convergence to the optimal QS-value under a set of minimal assumptions. Within this framework, the learning updates are restricted exclusively to the states present in the dataset. A rigorous definition of state reachability is indispensable for this setting:
@@ 4 State-constrained QS-learning
Definition 4.1 implies that a state is reachable if there exists an action that, when executed in the environment, leads to that state. This definition allows multiple reachable next states to be evaluated rather than a single state-action pair. By evaluating multiple reachable next states, state-constrained QS-learning allows for more flexibility in the learning updates compared to batch-constrained approaches, which are limited to evaluating only existing state-action pairs in the dataset.
4.1 Theory
In this section, we initially adapt the theorems presented in BCQL [22] to suit the state-constrained context (Theorems 4.4 - 4.6). All our subsequent theorems are based on QS-learning. However, it is important to note that these theorems still hold if QA-learning is used instead, with an inverse model defined to evaluate actions between states and reachable next states.
Our theory operates under the following assumptions: (A1) the environment is deterministic; (A2) the rewards are bounded such that $\forall(s,s^{\prime}),|r(s,s^{\prime})|\leq c$; (A3) the QS-values, $Q(s,s^{\prime})$, are initialised to finite values; and (A4) the discount factor is set such that $10\leq\gamma<$ 1.
First, we show that under these assumptions, QS-learning converges to the expert QS-value.
@@ 4.1 Theory
assuming each $(s,s^{\prime})$ pair is visited infinitely often, then as updates are performed, the agent’s hypothesis $Q_{n}(s,s^{\prime})$ after the $n$th update converges to the expert QS-value, $Q^{*}(s,s^{\prime})$, as $n\rightarrow\infty$ for all $s,s^{\prime}$.
@@
This follows from the convergence of QA-learning in a deterministic MDP [39]. For brevity and clarity, the full proof is shown in the Appendix. ∎
The convergence of QS-learning in a deterministic MDP, as shown in Theorem 4.2, is crucial for establishing the convergence of SCQL. Since SCQL is based on QS-learning and operates in a deterministic MDP, Theorem 4.2 provides the foundation for proving the convergence and optimality of SCQL under certain assumptions. We now demonstrate that learning the value function from the dataset $\mathcal{D}$ is equivalent to determining the value function of an associated MDP, denoted as $\mathcal{M}_{\mathcal{S}}$. Intuitively, $\mathcal{M}_{\mathcal{S}}$ is the MDP where all transitions are possible between reachable states found in $\mathcal{D}$.
@@
(State-constrained MDP) Let the state-constrained MDP $\mathcal{M}_{\mathcal{S}}=(\mathcal{S},\mathcal{A},\mathcal{P}_{\mathcal{S}},\mathcal{R},\gamma)$. Here, both $\mathcal{S}$ and $\mathcal{A}$ remain identical to those in the original MDP, $\mathcal{M}$. The transition probability is given by: $a=I(s,s^{\prime})\in\mathcal{A}$
@@
The reward function and discount factor remain the same as the original MDP.
The transition probability $p_{\mathcal{S}}(s^{\prime}|s,a)$ in the state-constrained MDP is defined such that transitions are possible only between states that are reachable from each other according to the state reachability definition. This ensures that the state-constrained MDP captures the essential dynamics of the original MDP while focusing on the states present in the dataset. The reward $r(s,s^{\prime})$ is assigned as the original reward defined in $\mathcal{M}$. Note that such states $s$ and $s^{\prime}$ may not exist as a pair $(s,s^{\prime})$ in $\mathcal{D}$. For the case where $s$ is absent from the dataset, the rewards are set to the initialised values $Q(s,s^{\prime})$. Importantly, the $s$ and $s^{\prime}$ in Definition 4.3 both exist in the dataset but not as a pair; this means that more transitions exist under this definition than in the batch-constrained MDP defined in [22].
@@
By sampling $s$ from the dataset $\mathcal{D}$, sampling $s^{\prime}$ from $\mathcal{SR}_{\mathcal{M}}(s)$ and performing QS-learning on all reachable state-next state pairs, QS-learning converges to the optimal value function of the state-constrained MDP $\mathcal{M}_{\mathcal{S}}$.
@@
Theorem 4.4 establishes that performing QS-learning on all reachable state-next state pairs from the dataset converges to the optimal value function of the state-constrained MDP. This result is crucial for understanding the convergence and optimality properties of SCQL, as it shows that QS-learning effectively learns the optimal value function of the state-constrained MDP, which is closely related to the original MDP. We are now ready to define the state-constrained QS-learning (SCQL) update which is similar to the BCQL formulation except now the maximisation is constrained to the states rather than state-action pairs in the dataset. This formulation allows the maximisation to be taken over more values composing more accurate Q-values while still staying close to the dataset:
$Q(s,s^{\prime})\leftarrow(1-\alpha)Q(s,s^{\prime})+\alpha\Bigl{[}r(s,s^{\prime})+\gamma\max_{\begin{subarray}{c}s^{\prime\prime}s.t.s^{\prime\prime}\in\mathcal{D}\\ s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})\end{subarray}}Q(s^{\prime},s^{\prime\prime})\Bigr{]}.$ (6)
SCQL, Eq. (6), converges under the identical conditions as traditional QS-learning, primarily because the state-constrained setting is non-limiting whenever every state in the MDP is observed.
@@
Under assumptions A1-4 and assuming every state $s$ is encountered infinitely, SCQL converges to the optimal value function, denoted as $Q^{*}$.
@@
$\pi_{s}^{*}(s)=\mathop{\arg\max}\limits_{\begin{subarray}{c}s^{\prime}~{}\mathrm{s.t.}s^{\prime}\in\mathcal{D}\\ s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)\end{subarray}}Q^{*}(s,s^{\prime}).$ (7)
Here, the maximisation is taken over next states from the dataset and that are reachable from the current state. We now demonstrate that Eq. (7) represents the optimal state-constrained policy.
@@
Under assumptions A1-4 and assuming every $s$ is visited infinitely, then SCQL converges to $Q^{\pi}_{\mathcal{S}}(s,s^{\prime})$, with the optimal state-constrained policy defined by Eq. (7) and $s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)$.
@@
In a deterministic MDP, SCQL will produce a policy that is never worse than the policy produced by BCQL.
@@
From Theorems 4.6 and 4.7, the state-constrained approach poses fewer restrictions than the batch-constrained counterpart while also always producing a policy at least as good. SCQL can perform more Q-value updates and maximize over more states than BCQL because it considers all reachable state-next state pairs from the dataset, rather than being limited to the explicit state-action pairs present in the dataset. This allows SCQL to exploit the structure of the state-constrained MDP more effectively and learn a better policy, with less required data. In the next section, this improvement is elaborated through an illustrative example which shows SCQL excelling even with a limited dataset.

### v2

# algorithmic

@@ Abstract
Traditional offline reinforcement learning (RL) methods predominantly operate in a batch-constrained setting. This confines the algorithms to a specific state-action distribution present in the dataset, reducing the effects of distributional shift but restricting the policy to seen actions. In this paper, we alleviate this limitation by introducing state-constrained offline RL, a novel framework that focuses solely on the dataset’s state distribution. This approach allows the policy to take high-quality out-of-distribution actions that lead to in-distribution states, significantly enhancing learning potential. The proposed setting not only broadens the learning horizon but also improves the ability to combine different trajectories from the dataset effectively, a desirable property inherent in offline RL. Our research is underpinned by theoretical findings that pave the way for subsequent advancements in this area. Additionally, we introduce StaCQ, a deep learning algorithm that achieves state-of-the-art performance on the D4RL benchmark datasets and aligns with our theoretical propositions. StaCQ establishes a strong baseline for forthcoming explorations in this domain.
@@ 1 Introduction
Offline RL aims to derive an optimal policy solely from a fixed dataset of pre-collected experiences, without further interactions with the environment (Lange et al., 2012; Levine et al., 2020). It prohibits online exploration, meaning that effective policies must be constructed solely from the evidence provided by an unknown, potentially sub-optimal behaviour policy active in the environment. This approach is especially suited for real-world scenarios where executing sub-optimal actions can be dangerous, time-consuming, or costly, yet there is an abundance of prior data. Relevant applications include robotics (Singh et al., 2020; Kumar et al., 2021; Sinha et al., 2022), long-term healthcare treatment plans (Tang & Wiens, 2021; Tang et al., 2022; Shiranthika et al., 2022), and autonomous driving (Shi et al., 2021; Fang et al., 2022; Diehl et al., 2023). Despite its practical appeal, offline RL faces a significant distributional shift challenge (Kumar et al., 2019). This phenomenon arises when attempting to estimate the values of actions not present in the dataset, often manifesting as an overestimation. As a result, out-of-distribution (OOD) actions are perceived as more valuable than they truly are, causing the agent to select sub-optimal actions and leading to the accumulation of errors (Fujimoto et al., 2019).
Currently available offline RL techniques mitigate the issue of distributional shift through two primary strategies. The first constrains the policy, during training, to take actions close to dataset actions (Fujimoto et al., 2019; Kumar et al., 2019; Wu et al., 2019; Siegel et al., 2020; Kostrikov et al., 2021a; Zhou et al., 2021; Fujimoto & Gu, 2021). The second revolves around conservatively estimating the value of OOD actions (Kumar et al., 2020; Yu et al., 2021; An et al., 2021). Both strategies aim to anchor the learning algorithm to the state-action pairs within the dataset, an approach termed batch-constrained (Fujimoto et al., 2019). By doing so, these methods seek to minimise the impact of distributional shift. However, adhering strictly to the state-action distribution can be restrictive when the optimal action is not found in the dataset. Thus, recent techniques seek a balance between exploiting potentially valuable OOD actions and minimising distributional shift repercussions.
@@
Constraining exclusively to the dataset’s state distribution could significantly reduce the requisite dataset size. Instead of needing all state-action pairs, the focus shifts to states alone. This approach allows the agent to initiate OOD actions, provided they lead to a known state, a form of safe action exploration which avoids distributional shift. Such flexibility can enhance the capability of offline RL algorithms in trajectory stitching — combining sub-optimal experiences into improved trajectories. Rather than connecting based on actions from different trajectories, the state-constrained approach facilitates stitching by leveraging adjacent states. This method relies on understanding these proximate states, a concept we label as reachability, which refers to the ability to reach a particular state from another state within the dataset.
Figure 1: Illustration of state reachability in a continuous state space. States ($s$) and next states ($s^{\prime}$) from the dataset are shown, with yellow circles representing additional reachable states. The state-constrained method identifies high-quality reachable states to mitigate distributional shift while improving policy performance.
This idea is illustrated in Figure 1, which shows the benefits of a state-constrained methodology. Most current offline RL methods constrain to the specific transition, $(s,a)$ or $(s,s^{\prime})$-pair, found in the dataset. Our state-constrained methodology can instead find alternative next states (yellow circles in Figure 1) and select the highest-value one. This approach effectively increases the number of transitions for the agent to learn from. It avoids distributional shift by updating on in-distribution states and leads to a higher-quality policy, as the policy is no longer restricted to explicitly observed actions.
In the state-constrained approach, understanding reachability is crucial for effectively stitching together sub-optimal trajectories and improving the overall policy. In some environments, reachability is easily delineated; for example, in grid-based environments like mazes, where the agent can move directly between adjacent cells. In other environments, reachability must be ascertained from the dataset, such as in complex robotic systems where state transitions depend on intricate dynamics.
@@
In this work, we make several key contributions. First, building on the concept of reachability, we introduce state-constrained offline RL. Unlike batch-constrained RL, which anchors learning to specific state–action pairs, state-constrained RL focuses on states, allowing more flexible action selection while still mitigating distributional shift. We also provide theoretical convergence guarantees under deterministic assumptions, showing that our framework yields policies whose actions are of higher or equal value compared to batch-constrained approaches. Another main contribution is StaCQ, a novel deep learning algorithm that learns a state-constrained value function and updates the policy to stay close to the highest-quality reachable states. We demonstrate competitive performance on multiple D4RL tasks, surpassing many state-of-the-art methods in locomotion and Antmaze. This positions StaCQ as a robust baseline for future work on state-constrained RL, similarly to how BCQ (Fujimoto et al., 2019) has served as a strong baseline for batch-constrained techniques, thereby advancing offline RL research.
@@ 2 Preliminaries
The RL framework involves an agent interacting with an environment, which is typically represented as a Markov Decision Process (MDP). Such an MDP is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively; $P=p(s^{\prime}|s,a)$ represents the environment transition dynamics; $R=r(s,a,s^{\prime})$ is the reward function for transitioning; and $10\leq\gamma<$ 1 is the discount factor (Sutton & Barto, 2018). As in Fujimoto et al. (2019), we focus on the deterministic MDP, where $p(s^{\prime}|s,a)=\{0,1\}$ and $r(s,a,s^{\prime})=r(s,s^{\prime})$. In RL, the agent’s objective is to identify an optimal policy, $\pi(s)$, that maximises the future discounted sum of rewards $\sum_{i=t}\gamma^{i-t}r_{i}(s_{i},s^{\prime}_{i})$.
QSA-values, $Q(s,a)=\mathbb{E}_{\pi}[\sum_{i=0}\gamma^{t}r_{t}(s_{t},a_{t})|s_{0}=s,a_{0}=a]$, are the expected sum of future discounted rewards for executing action $a$ in state $s$ and there after following policy $\pi$. Q-learning (Watkins & Dayan, 1992), denoted as QSA-learning in this paper, estimates the QSA-values under the assumption that future actions are chosen optimally,
@@
The optimal policy derived from QSA-learning identifies the best action that maximises the QSA-values, =$\pi^{*}(s)=\operatorname*{arg\,max}_{a}Q(s,a)$. In deterministic contexts, QSA-learning parallels QSS-learning (Edwards et al., 2020), which estimates the expected rewards upon transitioning from state $s$ to $s^{\prime}$, followed by optimal decisions:
@@
QSS-learning enables QSS-values to be learned for transitioning, without evaluating actions. The optimal policy via QSS-learning discerns the most advantageous subsequent state to maximise QSS-values, $\pi_{s}^{*}(s)=\operatorname*{arg\,max}_{s^{\prime}}Q(s,s^{\prime})$. An action can then be retrieved from an inverse dynamics model as $a=I(s,\pi_{s}^{*}(s))$, a strategy originally designed to address QSA-learning’s challenges in redundant action spaces, i.e. where multiple actions lead to the same next state. The methods presented in this paper build upon QSS-learning, using it to avoid evaluating OOD actions between reachable state pairs.
In offline RL, the agent aims to discover an optimal policy; however, it must solely rely on a fixed dataset without further interactions with the environment (Lange et al., 2012; Levine et al., 2020). The dataset comprises trajectories made up of transitions consisting of the current state, action, next state, and the reward for transitioning, where actions are chosen based on an unknown behaviour policy, $\pi_{\beta}$. Batch-Constrained QSA-learning (BCQL) (Fujimoto et al., 2019) adapts QSA-learning for the offline setting by restricting the optimisation to state-action pairs present in the dataset:
@@
This approach is restrictive since convergence to the optimal value function in all states is only ensured if every optimal $(s,a)$-pair from the MDP resides within the dataset. Such a limitation aims to sidestep extrapolation errors resulting from distributional shifts. Nonetheless, this framework places a significant constraint on any learning algorithm. Subsequent sections relax this constraint: instead of adhering strictly to state-action pairs (batch-constrained), the learning process is only bound by states (state-constrained). In the following section, we introduce the state-constrained framework and provide a new algorithm called state-constrained QSS-learning (SCQL). SCQL is based from QSS-learning which is useful to evaluate state and reachable next state pairs, where both states exist in the dataset, avoiding the distributional shift issue in offline RL. We show that, under minor assumptions, SCQL converges to the optimal QSS-value and produces a less-restrictive policy than BCQL.
3 State-constrained QSS-learning
In this section, we provide a formal introduction to state-constrained QSS-learning and establish its convergence to the optimal QSS-value under a set of minimal assumptions. Within this framework, the learning updates are restricted exclusively to the states present in the dataset. A rigorous definition of state reachability is indispensable for this setting:
@@ 3 State-constrained QSS-learning
Definition 3.1 implies that a state is reachable if there exists an action that, when executed in the environment, leads to that state. This definition allows multiple reachable next states to be evaluated rather than a single state-action pair. State-constrained QSS-learning enables more flexible learning updates by evaluating multiple reachable next states, as shown in Figure 1 where the agent can now learn from all reachable states rather than the single explicit next state. This flexibility leads to more robust policies that can better generalise beyond the specific transitions seen in the dataset.
3.1 Theoretical foundations
In this section, we initially adapt the theorems presented for BCQL (Fujimoto et al., 2019) to suit the state-constrained context (Theorems 3.4 - 3.6). All our subsequent theorems are proposed based on learning an optimal QSS-value. However, it is important to note that these theorems still hold if QSA-learning is used instead, with an inverse model defined to evaluate actions between states and reachable next states.
Our theory operates under the following assumptions: (A1) the environment is deterministic; (A2) the rewards are bounded such that $\forall(s,s^{\prime}),|r(s,s^{\prime})|\leq c$; (A3) the QSS-values, $Q(s,s^{\prime})$, are initialised to finite values; and (A4) the discount factor is set such that $10\leq\gamma<$ 1.
First, we show that under these assumptions, QSS-learning converges to the optimal QSS-value.
@@ 3.1 Theoretical foundations
assuming each $(s,s^{\prime})$ pair is visited infinitely often, let $Q_{n}(s,s^{\prime})$ be the value from the $n$th update of Eq. equation 4, then $Q_{n}(s,s^{\prime})$ converges to the optimal QSS-value, $Q^{*}(s,s^{\prime})$, as $n\rightarrow\infty$ for all $s,s^{\prime}$.
@@
This follows from the convergence of QSA-learning in a deterministic MDP (Mitchell, 1997). For brevity and clarity, the full proof is given in the Appendix. ∎
The convergence of QSS-learning in a deterministic MDP, as shown in Theorem 3.2, is crucial for establishing the convergence of SCQL. Since SCQL is based on QSS-learning and operates in a deterministic MDP, Theorem 3.2 provides the foundation for proving the convergence and optimality of SCQL under certain assumptions. We now demonstrate that learning the value function from the dataset $\mathcal{D}$ is equivalent to determining the value function of an associated MDP, denoted as $\mathcal{M}_{\mathcal{S}}$. Intuitively, $\mathcal{M}_{\mathcal{S}}$ is the MDP where all transitions are possible between reachable states found in $\mathcal{D}$.
@@
(State-constrained MDP) Let the state-constrained MDP $\mathcal{M}_{\mathcal{S}}=(\mathcal{S},\mathcal{A},\mathcal{P}_{\mathcal{S}},\mathcal{R},\gamma)$. Here, both $\mathcal{S}$ and $\mathcal{A}$ remain identical to those in the original MDP, $\mathcal{M}$ and $s_{\text{terminal}}$ is an additional terminating state. The transition probability is given by: $a=I(s,s^{\prime})\in\mathcal{A}$
@@
The reward function and discount factor remain the same as the original MDP. Except for the terminal state where $r(s,s_{\text{terminal}})$, is set to the initialised value of $Q(s,s^{\prime})$.
The transition probability $p_{\mathcal{S}}(s^{\prime}|s,a)$ in the state-constrained MDP is defined such that transitions are possible only between states that are reachable from one to the other according to the state reachability definition. This ensures that the state-constrained MDP captures the essential dynamics of the original MDP while focusing on the states present in the dataset. The reward $r(s,s^{\prime})$ is assigned as the original reward defined in $\mathcal{M}$. For the case where $s$ is absent from the dataset, the rewards are set to the initialised values $Q(s,s^{\prime})$. Importantly, the $s$ and $s^{\prime}$ in Definition 3.3 both exist in the dataset but may not exist as a pair $(s,s^{\prime})$; this means that more transitions exist under this definition than in the batch-constrained MDP defined in Fujimoto et al. (2019).
@@
By sampling $s$ from $\mathcal{D}$, sampling $s^{\prime}$ from $\mathcal{SR}_{\mathcal{M}}(s)$ and performing QSS-learning on all reachable state-next state pairs, QSS-learning converges to the optimal value function of the state-constrained MDP $\mathcal{M}_{\mathcal{S}}$.
@@
Theorem 3.4 establishes that performing QSS-learning on all reachable state-next state pairs from the dataset converges to the optimal value function of the state-constrained MDP. This result is crucial for understanding the convergence and optimality properties of SCQL, as it shows that QSS-learning effectively learns the optimal value function of the state-constrained MDP, which is closely related to the original MDP.
We are now ready to define the state-constrained QSS-learning (SCQL) update which is similar to the BCQL formulation except now the maximisation is constrained to the states rather than state-action pairs in the dataset. This formulation allows the maximisation to be taken over more values composing more accurate Q-values while still staying close to the dataset:
$Q(s,s^{\prime})\leftarrow(1-\alpha)Q(s,s^{\prime})+\alpha\Bigl{[}r(s,s^{\prime})+\gamma\max_{\begin{subarray}{c}s^{\prime\prime}s.t.s^{\prime\prime}\in\mathcal{D}\\ \cap\\ s^{\prime\prime}\in\mathcal{SR}_{\mathcal{M}}(s^{\prime})\end{subarray}}Q(s^{\prime},s^{\prime\prime})\Bigr{]}.$ (6)
SCQL, Eq. equation 6, converges under the identical conditions as traditional QSS-learning, primarily because the state-constrained setting is non-limiting whenever every state in the MDP is observed.
@@
Under assumptions A1-4 and assuming every state s is encountered infinitely, let $Q_{n}(s,s^{\prime})$ be the value from the $n$th update of Eq.equation 6, the update rule of SCQL, then $Q_{n}(s,s^{\prime})$ converges to the optimal QSS-value $Q^{*}(s,s^{\prime})$, as $n\rightarrow\infty$ for all $(s,s^{\prime})$.
@@
$\pi_{s}^{*}(s)=\operatorname*{arg\,max}_{\begin{subarray}{c}s^{\prime}~{}\mathrm{s.t.}s^{\prime}\in\mathcal{D}\\ \cap\\ s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)\end{subarray}}Q^{*}(s,s^{\prime}).$ (7)
Here, the maximisation is taken over next states from the dataset and that are reachable from the current state. We now demonstrate that Eq. equation 7 represents the optimal state-constrained policy.
@@
Under assumptions A1-4 and assuming every state s is encountered infinitely, let $Q_{n}(s,s^{\prime})$ be the value from the $n$th update of Eq.equation 6, the update rule of SCQL, then $Q_{n}(s,s^{\prime})$ converges to $Q^{\pi}_{\mathcal{S}}(s,s^{\prime})$, the optimal QSS-value computed from states from $\mathcal{D}$, with the optimal state-constrained policy defined by Eq. equation 7 where $s\in\mathcal{D}$ and $s^{\prime}\in\mathcal{SR}_{\mathcal{M}}(s)\cap\mathcal{D}$.
@@
Let $\pi_{\text{BCQL}}$ and $\pi_{\text{SCQL}}$ be the policies produced by BCQL and SCQL respectively. Then, in a deterministic MDP, $V_{\pi_{\text{SCQL}}}(s)\geq V_{\pi_{\text{BCQL}}}(s)$, $\forall s\in\mathcal{D}$.
@@
From Theorems 3.6 and 3.7, the state-constrained approach poses fewer restrictions than the batch-constrained counterpart while also always producing a policy at least as good. SCQL can perform more Q-value updates and maximise over more states than BCQL because it considers all reachable state-next state pairs from the dataset, rather than being limited to the explicit state-action pairs present in the dataset. This allows SCQL to exploit the structure of the state-constrained MDP more effectively and learn a better policy, with less required data. In the next section, this improvement is illustrated through an example which shows SCQL excelling even with a limited dataset.